# Rethinking Irregular Time Series Forecasting: A Simple yet Effective Baseline

### v2

@@ 1 Introduction
Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao et al., 2018; Brouwer et al., 2019; Kidger et al., 2020; Ham et al., 2019; Vio et al., 2013). Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making. However, the inherent irregularity of observations and missing data (Yalavarthi et al., 2024; Zhang et al., 2024) in IMTS pose significant challenges to IMTSF modeling.
To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), have adopted the Fixed Patching approach—see Figure 1a. This approach divides the time series into fixed-length patches with equal intervals among them. However, this approach has notable limitations: 1) Uneven Information Density Across Patches: Fixed Patching struggles to adapt to local variations in data density, leading to uneven information density among patches. For example, sparse patches (where the number of observations is limited) may result in insufficient feature extraction, yet dense patches (where the number of observations is abundant) may contain redundant information or noise, thereby impairing the extracted features. 2) Inappropriate Segmentation of Key Semantic Information: Fixed Patching risks splitting critical dynamic information, which hampers the model’s ability to capture the complete semantic context. Therefore, the first challenge is to design an adaptive patching approach that can effectively adapt to local information density variations in IMTS and capture the complete semantic context.
@@
Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times. For example, neural ODEs-based models (Chen et al., 2018; Rubanova et al., 2019; Gravina et al., 2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin, 2021; Chen et al., 2018). GNNs-based models (Yalavarthi et al., 2024; Zhang et al., 2024) are hindered by the overhead of complex graph construction and multi-round node information aggregation (Wu et al., 2021); Transformer-based models (Chen et al., 2023; Zhang et al., 2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al., 2024). These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS. While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios. Notably, in regular time series forecasting task, models such as DLinear (Zeng et al., 2023), SparseTSF (Lin et al., 2024b), and CycleNet (Lin et al., 2024a) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters. Therefore, the second challenge is to design a model for IMTSF that is both highly accurate and computationally efficient.
@@
  • We conduct extensive experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency. Additionally, all datasets and code are available at https://anonymous.4open.science/r/APN-23D0.
@@ 2.1 Irregular Multivariate Time Series Forecasting
IMTSF is crucial for key domains such as healthcare and climate science. The inherent characteristics of IMTS, such as non-uniform sampling intervals and asynchronous channels, present significant challenges to regular time series forecasting models. To address these characteristics, researchers have proposed various IMTSF models. Some models employ approaches based on continuous-time dynamics (Chen et al., 2018; Rubanova et al., 2019; Schirmer et al., 2022; Brouwer et al., 2019), utilizing ordinary or stochastic differential equations (ODE/SDE) to adapt to irregular sampling points. For instance, Neural Flows (Bilos et al., 2021) proposes directly modeling the solution curves of ODEs, thereby avoiding the costly numerical integration steps in traditional ODE solvers. GRU-ODE-Bayes (Brouwer et al., 2019) innovatively combines the idea of Gated Recurrent Units (GRU) with ODEs and effectively handles sparse observational data through a Bayesian update mechanism. Other models leverage graph neural networks and attention mechanisms (Yalavarthi et al., 2024; Zhang et al., 2024; Liu et al., 2025) to capture complex dependencies in IMTS. For example, GraFITi (Yalavarthi et al., 2024) transforms IMTS into sparse bipartite graphs and predicts edge weights through GNNs. T-PatchGNN (Zhang et al., 2024) innovatively segments irregular sequences into time-aligned patches and combines Transformer and adaptive GNNs to handle intra-patch and inter-patch dependencies, respectively.
@@ 2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
The patch-based strategy, which has proven successful for regular time series (Nie et al., 2023; Wang et al., 2022), has been adapted for IMTS Forecasting (IMTSF). Initial adaptations, such as T-PatchGNN (Zhang et al., 2024) and TimeCHEAT (Liu et al., 2025), employed a fixed-span patching strategy. However, this rigid segmentation is ill-suited for the non-uniform data distribution of IMTS. It often creates patches with highly variable information content—some being too sparse for meaningful feature extraction, while others may be overly dense with redundant data.
To overcome this, adaptive patching emerged as a natural evolution. One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (Huang et al., 2024)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling. While effective for dense, regular data, this methodology is fundamentally flawed for IMTS. Interpolating across sparse or missing intervals can fabricate misleading data points, introducing significant artifacts and undermining model reliability.
@@ 3.3 Time-Aware Patch Aggregation (TAPA)
3.3.1 Adaptive Patching: Learning Dynamic Temporal Windows
The first step is to define the temporal scope of each patch. Instead of imposing pre-defined, rigid boundaries, TAPA learns a dynamic temporal window, $[t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$. This dynamism is the core mechanism by which TAPA adapts to the non-uniform distribution of data. For instance, in a sparse but critical region, the model can learn to expand a window’s width to ensure sparse yet vital signals are encompassed. Conversely, in a dense region with redundant data, a window can be strategically narrowed to focus on the most salient signals, thus preventing dilution from less informative points.
@@ 3.3.1 Adaptive Patching: Learning Dynamic Temporal Windows
3.3.2 Weighted Aggregation: From Raw Observations to Rich Representations
@@ 3.4 Query-based Weighted Aggregation
The resulting summary representations for all channels, $[H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$. A layer normalization is then applied to stabilize the input for the decoder. This lightweight aggregation mechanism is a direct testament to our design philosophy: by front-loading the complexity into TAPA, the subsequent modules can be elegantly simple yet powerful.
@@ 4.1 Setup
Baselines: To comprehensively evaluate the performance of APN, we compare it with nineteen baseline models. These baseline models can be broadly categorized into the following groups: (1) Regular MTSF models: including DLinear (Zeng et al., 2023), TimesNet (Wu et al., 2023), PatchTST (Nie et al., 2023), Crossformer (Wang et al., 2022), Graph Wavenet (Wu et al., 2019), MTGNN (Wu et al., 2020), StemGNN (Cao et al., 2020), CrossGNN (Huang et al., 2023), and FourierGNN (Yi et al., 2023). (2) IMTS classification/imputation models: including GRU-D (Che et al., 2018), SeFT (Horn et al., 2020), RainDrop (Zhang et al., 2022), Warpformer (Zhang et al., 2023), and mTAND (Shukla and Marlin, 2021). (3) IMTSF models: including Latent ODEs (Rubanova et al., 2019), CRU (Schirmer et al., 2022), and Neural Flows (Bilos et al., 2021), T-PATCHGNN (Zhang et al., 2024), and TimeCHEAT (Liu et al., 2025). For detailed information on the baseline models, please refer to Appendix A.2.
@@ 4.5 Scalability and Efficiency Analysis
To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (TimeCHEAT (Liu et al., 2025), CRU (Schirmer et al., 2022), T-PatchGNN (Zhang et al., 2024)) on the PhysioNet dataset, with the batch size uniformly set to 32—see Figure 4. The results demonstrate that APN exhibits significant advantages across all key efficiency metrics. This is attributed to its core TAPA module, which efficiently generates information-condensed and regularized initial representations. These representations support a highly streamlined information aggregation and prediction architecture, ultimately achieving a lightweight and highly efficient model design.

### v3

@@ 1 Introduction
Irregular Multivariate Time Series (IMTS) data are widely observed in various domains such as healthcare, biomechanics, climate science, and astronomy (Yao, Bi, and Chen 2018; Brouwer et al. 2019; Kidger et al. 2020; Ham, Kim, and Luo 2019; Vio, Díaz-Trigo, and Andreani 2013). Irregular Multivariate Time Series Forecasting (IMTSF) is a crucial research task that provides valuable insights for early warning and proactive decision-making. However, the inherent irregularity of observations and missing data (Yalavarthi et al. 2024; Zhang et al. 2024) in IMTS pose significant challenges to IMTSF modeling.
To address this challenge, recent IMTSF methods, such as T-PatchGNN (Zhang et al. 2024) and TimeCHEAT (Liu, Cao, and Chen 2025), have adopted the Fixed Patching approach—see Figure 1a. This approach divides the time series into fixed-length patches with equal intervals among them. However, this approach has notable limitations: 1) Uneven Information Density Across Patches: Fixed Patching struggles to adapt to local variations in data density, leading to uneven information density among patches. For example, sparse patches (where the number of observations is limited) may result in insufficient feature extraction, yet dense patches (where the number of observations is abundant) may contain redundant information or noise, thereby impairing the extracted features. 2) Inappropriate Segmentation of Key Semantic Information: Fixed Patching risks splitting critical dynamic information, which hampers the model’s ability to capture the complete semantic context. Therefore, the first challenge is to design an adaptive patching approach that can effectively adapt to local information density variations in IMTS and capture the complete semantic context.
@@
Meanwhile, existing IMTSF models generally suffer from high computational costs and long running times. For example, neural ODEs-based models (Chen et al. 2018; Rubanova, Chen, and Duvenaud 2019; Gravina et al. 2024) require computationally intensive numerical solvers to accurately model continuous-time dynamics (Shukla and Marlin 2021; Chen et al. 2018). GNNs-based models (Yalavarthi et al. 2024; Zhang et al. 2024) are hindered by the overhead of complex graph construction and multi-round node information aggregation (Wu et al. 2021); Transformer-based models (Chen et al. 2023; Zhang et al. 2023), which utilize multi-layer self-attention mechanisms and feedforward networks, often result in large parameter scales (Kim et al. 2024). These models typically construct computationally intensive and parameter-heavy complex architectures to effectively handle the intricate dependencies and dynamic changes in IMTS. While this design enhances the model’s performance to some extent, it also incurs high computational costs and long runtimes, limiting its practical application in resource-constrained scenarios. Notably, in regular time series forecasting task, models such as DLinear (Zeng et al. 2023), SparseTSF (Lin et al. 2024b), and CycleNet (Lin et al. 2024a) have significantly reduced computational overhead and achieved competitive prediction accuracy by adopting simple architectures with fewer parameters. Therefore, the second challenge is to design a model for IMTSF that is both highly accurate and computationally efficient.
@@
  • We conduct extensive experiments on multiple datasets. The results show that APN outperforms existing SOTA baselines in both forecasting accuracy and computational efficiency.
@@ 2.1 Irregular Multivariate Time Series Forecasting
IMTSF is crucial for key domains such as healthcare and climate science. The inherent characteristics of IMTS, such as non-uniform sampling intervals and asynchronous channels, present significant challenges to regular time series forecasting models. To address these characteristics, researchers have proposed various IMTSF models. Some models employ approaches based on continuous-time dynamics (Chen et al. 2018; Rubanova, Chen, and Duvenaud 2019; Schirmer et al. 2022; Brouwer et al. 2019), utilizing ordinary or stochastic differential equations (ODE/SDE) to adapt to irregular sampling points. For instance, Neural Flows (Bilos et al. 2021) proposes directly modeling the solution curves of ODEs, thereby avoiding the costly numerical integration steps in traditional ODE solvers. GRU-ODE-Bayes (Brouwer et al. 2019) innovatively combines the idea of Gated Recurrent Units (GRU) with ODEs and effectively handles sparse observational data through a Bayesian update mechanism. Other models leverage graph neural networks and attention mechanisms (Yalavarthi et al. 2024; Zhang et al. 2024; Liu, Cao, and Chen 2025) to capture complex dependencies in IMTS. For example, GraFITi (Yalavarthi et al. 2024) transforms IMTS into sparse bipartite graphs and predicts edge weights through GNNs. T-PatchGNN (Zhang et al. 2024) innovatively segments irregular sequences into time-aligned patches and combines Transformer and adaptive GNNs to handle intra-patch and inter-patch dependencies, respectively.
@@ 2.2 Progress in Patch-based Irregular Multivariate Time Series Forecasting
The patch-based strategy, which has proven successful for regular time series (Nie et al. 2023; Wang et al. 2022), has been adapted for IMTS Forecasting (IMTSF). Initial adaptations, such as T-PatchGNN (Zhang et al. 2024) and TimeCHEAT (Liu, Cao, and Chen 2025), employed a fixed-span patching strategy. However, this rigid segmentation is ill-suited for the non-uniform data distribution of IMTS. It often creates patches with highly variable information content—some being too sparse for meaningful feature extraction, while others may be overly dense with redundant data.
To overcome this, adaptive patching emerged as a natural evolution. One prominent approach, primarily explored in the context of regular time series (e.g., HDMixer (Huang et al. 2024)), involves learning flexible patch boundaries and then generating regularized patch representations through interpolation-based resampling. While effective for dense, regular data, this methodology is fundamentally flawed for IMTS. Interpolating across sparse or missing intervals can fabricate misleading data points, introducing significant artifacts and undermining model reliability.
@@ 3.3 Time-Aware Patch Aggregation (TAPA)
Adaptive Patching: Learning Dynamic Temporal Windows
The first step is to define the temporal scope of each patch. Instead of imposing pre-defined, rigid boundaries, TAPA learns a dynamic temporal window, $t_{p}^{left,n},t_{p}^{right,n}]$, for each target patch $p$. This dynamism is the core mechanism by which TAPA adapts to the non-uniform distribution of data. For instance, in a sparse but critical region, the model can learn to expand a window’s width to ensure sparse yet vital signals are encompassed. Conversely, in a dense region with redundant data, a window can be strategically narrowed to focus on the most salient signals, thus preventing dilution from less informative points.
@@
Weighted Aggregation: From Raw Observations to Rich Representations
@@ 3.4 Query-based Weighted Aggregation
The resulting summary representations for all channels, $H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$. A layer normalization is then applied to stabilize the input for the decoder. This lightweight aggregation mechanism is a direct testament to our design philosophy: by front-loading the complexity into TAPA, the subsequent modules can be elegantly simple yet powerful.
@@ 4.1 Setup
Baselines: To comprehensively evaluate the performance of APN, we compare it with nineteen baseline models. These baseline models can be broadly categorized into the following groups: (1) Regular MTSF models: including DLinear (Zeng et al. 2023), TimesNet (Wu et al. 2023), PatchTST (Nie et al. 2023), Crossformer (Wang et al. 2022), Graph Wavenet (Wu et al. 2019), MTGNN (Wu et al. 2020), StemGNN (Cao et al. 2020), CrossGNN (Huang et al. 2023), and FourierGNN (Yi et al. 2023). (2) IMTS classification/imputation models: including GRU-D (Che et al. 2018), SeFT (Horn et al. 2020), RainDrop (Zhang et al. 2022), Warpformer (Zhang et al. 2023), and mTAND (Shukla and Marlin 2021). (3) IMTSF models: including Latent ODEs (Rubanova, Chen, and Duvenaud 2019), CRU (Schirmer et al. 2022), and Neural Flows (Bilos et al. 2021), T-PATCHGNN (Zhang et al. 2024), and TimeCHEAT (Liu, Cao, and Chen 2025). For detailed information on the baseline models, please refer to Appendix A.2.
@@ 4.5 Scalability and Efficiency Analysis
To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (TimeCHEAT (Liu, Cao, and Chen 2025), CRU (Schirmer et al. 2022), T-PatchGNN (Zhang et al. 2024)) on the PhysioNet dataset, with the batch size uniformly set to 32—see Figure 4. The results demonstrate that APN exhibits significant advantages across all key efficiency metrics. This is attributed to its core TAPA module, which efficiently generates information-condensed and regularized initial representations. These representations support a highly streamlined information aggregation and prediction architecture, ultimately achieving a lightweight and highly efficient model design.